# SCORE: Story Coherence and Retrieval Enhancement for AI Narratives

### v4

@@ Abstract
**footnotetext: Equal contribution.
@@ B-B1 Preliminaries
The details and underlying proof are provided in the section B-B.
@@
From lemma B.1, by constructing the Lagrangian
@@
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Score: 3
Evaluation Reasoning:
**Evaluation of Current Episode (26):**
1. **Character Consistency:**
- **School President BF:** His actions in this episode align with his established traits of being possessive and manipulative. His internal struggle between desire and restraint is consistent with his previous behavior of trying to balance his affection for Wakana with his controlling nature.
- **Wakana:** Her reactions are mostly consistent with her character, showing naivety and a desire to see the best in others. However, there is a discrepancy with the musical note hairpin-she fidgets with it despite it being lost in previous episodes, which is a continuity error.
- **Mei:** Meis concern about Wakanas safety and her suspicion towards the School President BF align with her established role as a protective friend.
2. **Plot Progression:**
- This episode contributes to the story by escalating the tension between the School President BF and the other characters. The suggestion of living together pushes the narrative of the BFs increasing control over Wakana. Meis intervention hints at potential conflict, which adds intrigue and sets the stage for future developments.
3. **Continuity and Consistency in Story Elements:**
- There is a significant continuity error with the appearance of the musical note hairpin, which was previously lost. This oversight impacts the narrative coherence, as the hairpins loss was a symbol of Wakanas emotional turmoil. Additionally, Wakanas phone being manipulated by the School President BF without her knowledge adds to her vulnerability, fitting the narrative but needs clearer exposition.
**Score: 3 (Fair)**
**Justification:**
The episode maintains character consistency and advances the plot effectively, with realistic emotional responses from the characters. However, noticeable continuity errors, particularly with the musical note hairpin, and a slight lack of clarity regarding Wakanas phone manipulation, detract from the overall coherence. These issues affect the narrative flow and reduce the episodes effectiveness, resulting in a score of 3.
@@
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Character Consistency:
In Episode 26, the actions and dialogues of the main characters remain mostly consistent with their established traits. The School President BF continues to exhibit his possessive, manipulative tendencies, which align with his characters controlling nature. His dialogue remains calculating and laced with undertones of jealousy, particularly when interacting with Kenji. The School President BFs obsessive behavior and desire to maintain control over Wakana are still central to his character, and his actions reflect this consistent personality trait.
Wakana maintains her characteristic obliviousness to the undercurrents of jealousy, which is also in line with her established naivety. However, theres a subtle evolution in her character as she shows more emotional depth and awareness of her surroundings, suggesting growth. Kenji, as always, remains timid and caught between his emotions for Wakana and fear of School President BF.
The characters stay consistent overall, but there is slight development, particularly with Wakana, making the episode believable within the context of the storyline.
Plot Progression:
This episode contributes significantly to the plot, moving the story forward by deepening the tensions between the School President BF and Kenji. The confrontation between the two characters intensifies, and the stakes are raised with the School President BF exerting more overt pressure on Kenji to leave the school. The introduction of Kenjis potential transfer provides a dramatic twist, which escalates the conflict and sets up further developments.
The episode effectively builds suspense, especially with the School President BFs plans to force Kenji to leave, creating anticipation for future episodes. The plot thread involving Wakanas emotional conflict adds another layer of complexity to the narrative, keeping the audience engaged. The pacing is good, and the elements introduced feel like logical progressions of the ongoing storylines.
Continuity and Consistency in Story Elements:
There are no major continuity errors in this episode. The story elements introduced are consistent with previous episodes, and there are no objects or events that contradict prior established facts. The ongoing conflict between the School President BF and Kenji, as well as Wakanas central role, remains intact without any glaring contradictions or inconsistencies.
One minor point that could have been addressed more clearly is Wakanas hairpin, which has been a symbolic accessory throughout the series. While it appears briefly, its importance to the emotional tone of the episode could have been emphasized further, as it has been a key item in previous episodes. However, this does not significantly impact the narrative coherence.
Score: 4 (Good)
The episode is mostly coherent with strong character consistency, good plot progression, and no significant continuity issues. The character dynamics evolve slightly, and the plot advances in a logical way, keeping the audience engaged. There are only minor areas where the narrative could have been enhanced, particularly with regards to the emotional depth of certain symbolic elements.
@@
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Score: 3
Evaluation Reasoning:
**Critical Evaluation of Episode 26:**
1. **Character Consistency:**
- The School President BFs actions and dialogue are consistent with his established traits of possessiveness and desire for Wakana, as seen in previous episodes. His internal struggle between desire and restraint aligns with his past behavior, showing his ongoing conflict.
- Wakanas innocence and naivety are also consistent, as she remains oblivious to the potential manipulation from the School President BF. However, there is an inconsistency with the mention of her musical note hairpin since it was previously noted as lost on a cliff. This could be an error unless it symbolizes her longing for the past.
2. **Plot Progression:**
- The scene contributes to the ongoing tension between Wakana and the School President BF while introducing the idea of moving in together, which could significantly alter their relationship dynamics.
- Meis concern about Wakana staying at the School President BFs place introduces a potential conflict and hints at the School Presidents manipulative behavior, maintaining suspense.
- The episode advances the plot by highlighting the growing control of the School President BF over Wakana, evidenced by his manipulation of her phone, which increases her dependency on him.
3. **Continuity and Consistency in Story Elements:**
- The episode has a notable inconsistency with the presence of the musical note hairpin, which was previously stated to be lost. This inconsistency undermines the narrative coherence.
- The manipulation of Wakanas phone is consistent with the School President BFs controlling behavior, but it needs further exploration to understand its implications fully.
**Score: 3 (Fair)**
**Justification:**
The episode is generally coherent with consistent character portrayals and logical plot progression. However, the inconsistency regarding the musical note hairpin affects the narrative flow and raises questions about continuity. Emotional responses are believable, but the episode could benefit from more clarity regarding the implications of key items, such as the manipulated phone and the lost hairpin, to enhance narrative coherence.
@@
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Evaluation of Episode 26
Character Consistency: In this episode, the School President BF continues to show his established traits of jealousy, possessiveness, and manipulation, which are consistent with his earlier portrayal. His actions towards Kenji, including his attempts to remove him as a threat to his relationship with Wakana, align with the yandere-like characterization presented in earlier episodes. Wakana, similarly, remains consistent in her obliviousness to the deeper conflict, maintaining her trusting and carefree personality, especially towards Kenji. There are no major inconsistencies in their behavior, and the progression of their personalities fits with the established dynamics of the story. However, Wakanas deep emotional response towards Kenji in this episode does suggest some growth or realization of his importance to her, which could potentially signal a shift in her character arc.
Plot Progression: This episode significantly advances the plot by heightening the tension between the School President BF and Kenji. The stakes are raised as the President overtly threatens Kenji, marking a clear escalation in his possessiveness and willingness to act on his jealousy. Kenjis reluctance to involve Wakana, despite being under intense pressure, adds complexity to the narrative, and the revelation that Kenji may be forced to leave the school contributes to the overall conflict. Wakanas emotional response hints at potential future conflict between her and the School President, as she may begin to question his actions and motivations. This scene effectively moves the story forward, setting the stage for a climax in their triangular relationship.
Continuity and Consistency in Story Elements: The episode maintains strong continuity with prior episodes, especially regarding key objects like Wakanas musical note hairpin, which has been a recurring symbol of her character. There are no glaring continuity errors, such as items reappearing without explanation. However, the escalating threat from the School President towards Kenji is beginning to stretch the believability of their relationship with Wakana, as her ongoing obliviousness to the intensity of the conflict between the two feels slightly contrived at this point. Additionally, the Presidents use of surveillance technology, though previously established, could benefit from further explanation or grounding in the storys logic to maintain narrative coherence.
Score: 4 (Good)
This episode is well-executed, with strong character consistency and meaningful plot advancement. The only drawback is Wakanas continued naivet\’{e}, which feels increasingly unrealistic given the circumstances. There is also room for more explicit explanation of some story elements, such as the Presidents use of surveillance, to avoid straining the narratives credibility.

### v5

@@ Abstract
**footnotetext: Equal contribution.footnotetext: Corresponding author.
@@ B-B1 Preliminaries
The details and underlying proof are provided in the section˜B-B.
@@
From lemma˜B.1, by constructing the Lagrangian
@@
Score: 3 Evaluation Reasoning: **Evaluation of Current Episode (26):** 1. **Character Consistency:** - **School President BF:** His actions in this episode align with his established traits of being possessive and manipulative. His internal struggle between desire and restraint is consistent with his previous behavior of trying to balance his affection for Wakana with his controlling nature. - **Wakana:** Her reactions are mostly consistent with her character, showing naivety and a desire to see the best in others. However, there is a discrepancy with the musical note hairpin-she fidgets with it despite it being lost in previous episodes, which is a continuity error. - **Mei:** Meis concern about Wakanas safety and her suspicion towards the School President BF align with her established role as a protective friend. 2. **Plot Progression:** - This episode contributes to the story by escalating the tension between the School President BF and the other characters. The suggestion of living together pushes the narrative of the BFs increasing control over Wakana. Meis intervention hints at potential conflict, which adds intrigue and sets the stage for future developments. 3. **Continuity and Consistency in Story Elements:** - There is a significant continuity error with the appearance of the musical note hairpin, which was previously lost. This oversight impacts the narrative coherence, as the hairpins loss was a symbol of Wakanas emotional turmoil. Additionally, Wakanas phone being manipulated by the School President BF without her knowledge adds to her vulnerability, fitting the narrative but needs clearer exposition. **Score: 3 (Fair)** **Justification:** The episode maintains character consistency and advances the plot effectively, with realistic emotional responses from the characters. However, noticeable continuity errors, particularly with the musical note hairpin, and a slight lack of clarity regarding Wakanas phone manipulation, detract from the overall coherence. These issues affect the narrative flow and reduce the episodes effectiveness, resulting in a score of 3.
@@
Character Consistency: In Episode 26, the actions and dialogues of the main characters remain mostly consistent with their established traits. The School President BF continues to exhibit his possessive, manipulative tendencies, which align with his characters controlling nature. His dialogue remains calculating and laced with undertones of jealousy, particularly when interacting with Kenji. The School President BFs obsessive behavior and desire to maintain control over Wakana are still central to his character, and his actions reflect this consistent personality trait. Wakana maintains her characteristic obliviousness to the undercurrents of jealousy, which is also in line with her established naivety. However, theres a subtle evolution in her character as she shows more emotional depth and awareness of her surroundings, suggesting growth. Kenji, as always, remains timid and caught between his emotions for Wakana and fear of School President BF. The characters stay consistent overall, but there is slight development, particularly with Wakana, making the episode believable within the context of the storyline. Plot Progression: This episode contributes significantly to the plot, moving the story forward by deepening the tensions between the School President BF and Kenji. The confrontation between the two characters intensifies, and the stakes are raised with the School President BF exerting more overt pressure on Kenji to leave the school. The introduction of Kenjis potential transfer provides a dramatic twist, which escalates the conflict and sets up further developments. The episode effectively builds suspense, especially with the School President BFs plans to force Kenji to leave, creating anticipation for future episodes. The plot thread involving Wakanas emotional conflict adds another layer of complexity to the narrative, keeping the audience engaged. The pacing is good, and the elements introduced feel like logical progressions of the ongoing storylines. Continuity and Consistency in Story Elements: There are no major continuity errors in this episode. The story elements introduced are consistent with previous episodes, and there are no objects or events that contradict prior established facts. The ongoing conflict between the School President BF and Kenji, as well as Wakanas central role, remains intact without any glaring contradictions or inconsistencies. One minor point that could have been addressed more clearly is Wakanas hairpin, which has been a symbolic accessory throughout the series. While it appears briefly, its importance to the emotional tone of the episode could have been emphasized further, as it has been a key item in previous episodes. However, this does not significantly impact the narrative coherence. Score: 4 (Good) The episode is mostly coherent with strong character consistency, good plot progression, and no significant continuity issues. The character dynamics evolve slightly, and the plot advances in a logical way, keeping the audience engaged. There are only minor areas where the narrative could have been enhanced, particularly with regards to the emotional depth of certain symbolic elements.
@@
Score: 3 Evaluation Reasoning: **Critical Evaluation of Episode 26:** 1. **Character Consistency:** - The School President BFs actions and dialogue are consistent with his established traits of possessiveness and desire for Wakana, as seen in previous episodes. His internal struggle between desire and restraint aligns with his past behavior, showing his ongoing conflict. - Wakanas innocence and naivety are also consistent, as she remains oblivious to the potential manipulation from the School President BF. However, there is an inconsistency with the mention of her musical note hairpin since it was previously noted as lost on a cliff. This could be an error unless it symbolizes her longing for the past. 2. **Plot Progression:** - The scene contributes to the ongoing tension between Wakana and the School President BF while introducing the idea of moving in together, which could significantly alter their relationship dynamics. - Meis concern about Wakana staying at the School President BFs place introduces a potential conflict and hints at the School Presidents manipulative behavior, maintaining suspense. - The episode advances the plot by highlighting the growing control of the School President BF over Wakana, evidenced by his manipulation of her phone, which increases her dependency on him. 3. **Continuity and Consistency in Story Elements:** - The episode has a notable inconsistency with the presence of the musical note hairpin, which was previously stated to be lost. This inconsistency undermines the narrative coherence. - The manipulation of Wakanas phone is consistent with the School President BFs controlling behavior, but it needs further exploration to understand its implications fully. **Score: 3 (Fair)** **Justification:** The episode is generally coherent with consistent character portrayals and logical plot progression. However, the inconsistency regarding the musical note hairpin affects the narrative flow and raises questions about continuity. Emotional responses are believable, but the episode could benefit from more clarity regarding the implications of key items, such as the manipulated phone and the lost hairpin, to enhance narrative coherence.
@@
Evaluation of Episode 26 Character Consistency: In this episode, the School President BF continues to show his established traits of jealousy, possessiveness, and manipulation, which are consistent with his earlier portrayal. His actions towards Kenji, including his attempts to remove him as a threat to his relationship with Wakana, align with the yandere-like characterization presented in earlier episodes. Wakana, similarly, remains consistent in her obliviousness to the deeper conflict, maintaining her trusting and carefree personality, especially towards Kenji. There are no major inconsistencies in their behavior, and the progression of their personalities fits with the established dynamics of the story. However, Wakanas deep emotional response towards Kenji in this episode does suggest some growth or realization of his importance to her, which could potentially signal a shift in her character arc. Plot Progression: This episode significantly advances the plot by heightening the tension between the School President BF and Kenji. The stakes are raised as the President overtly threatens Kenji, marking a clear escalation in his possessiveness and willingness to act on his jealousy. Kenjis reluctance to involve Wakana, despite being under intense pressure, adds complexity to the narrative, and the revelation that Kenji may be forced to leave the school contributes to the overall conflict. Wakanas emotional response hints at potential future conflict between her and the School President, as she may begin to question his actions and motivations. This scene effectively moves the story forward, setting the stage for a climax in their triangular relationship. Continuity and Consistency in Story Elements: The episode maintains strong continuity with prior episodes, especially regarding key objects like Wakanas musical note hairpin, which has been a recurring symbol of her character. There are no glaring continuity errors, such as items reappearing without explanation. However, the escalating threat from the School President towards Kenji is beginning to stretch the believability of their relationship with Wakana, as her ongoing obliviousness to the intensity of the conflict between the two feels slightly contrived at this point. Additionally, the Presidents use of surveillance technology, though previously established, could benefit from further explanation or grounding in the storys logic to maintain narrative coherence. Score: 4 (Good) This episode is well-executed, with strong character consistency and meaningful plot advancement. The only drawback is Wakanas continued naivet\’{e}, which feels increasingly unrealistic given the circumstances. There is also room for more explicit explanation of some story elements, such as the Presidents use of surveillance, to avoid straining the narratives credibility.